Qn 3. Consider a sequence {an}∞

n=0 of real numbers defined as follows:

a0 = 3, a1 = 2, a2 = 12, and

2an+3 − an+2 − 8an+1 + 4an = 0 for all n ≥ 0.

(i) What is a99?

(ii) Find a formula for 2a2n+1 − a2n in terms of n


1
Expert's answer
2022-02-28T14:06:59-0500

The characteristic polynomial of this recurrence relation is


"2r^3-r^2-8r+4=0"

"(r-2)(2r^2+3r-2)=0"

"(r-2)(r+2)(2r-1)=0"

The characteristic roots are "r_1=-2, r_2=1\/2, r_3=2."

Hence, the solutions to this recurrence relation are of the form


"a_n=\alpha_1(-2)^n+\alpha_2(\dfrac{1}{2})^n+\alpha_3(2)^n"

To find the constants "\alpha_1,\alpha_2,\alpha_3," use the initial conditions. This gives


"a_0=\alpha_1+\alpha_2+\alpha_3=3"

"a_1=-2\alpha_1+\dfrac{1}{2}\alpha_2+2\alpha_3=2"

"a_2=4\alpha_1+\dfrac{1}{4}\alpha_2+4\alpha_3=12"

"\alpha_1=1, \alpha_2=0, \alpha_3=2"

Hence, the unique solution to this recurrence relation and the given initial

conditions is the sequence {an} with


"a_n=(-2)^n+2(2)^n"

(i)


"a_{99}=(-2)^{99}+2(2)^{99}=(2)^{99}"

(ii)


"2a_{2n+1}-a_{2n}=2(-2)^{2n+1}+4(2)^{2n+1}"

"-((-2)^{2n}+2(2)^{2n})"

"=-4(2)^{2n}+8(2)^{2n}-(2)^{2n}-2(2)^{2n}"

"=(2)^{2n}=(4)^{n}"